Parth (parthkohli):

Challenge #2! Let \(a,b,c \rm ~ be ~ distinct~ numbers ~\in R\). The expressions given below are positive for all real \(x\) values. \(ax^2 + bx + c\) \(bx^2 + cx + a\) \(cx^2 + ax + b\) Let\[\alpha = \dfrac{ab + bc + ca}{a^2 + b^2 + c^2}\]Then: A. \(\alpha > 1 \) B. \(\alpha < 1\) C. \(\alpha > 1/4\) D. \(\alpha < 4\) Check ALL THAT APPLY.

ganeshie8 (ganeshie8):

maybe i would start here : Since the quadratics stay positive always, D < 0 b^2 -4ac < 0 c^2 - 4ab < 0 a^2 - 4bc < 0

Parth (parthkohli):

Right way, @ganeshie8.

ganeshie8 (ganeshie8):

Adding : b^2 -4ac < 0 c^2 - 4ab < 0 a^2 - 4bc < 0 --------------------- a^2 + b^2 + c^2 - 4(ab+bc+ca) < 0

mathslover (mathslover):

Or : \(a^2 + b^2 + c^2 < 4(ab + bc + ca) \) It would be easy now to conclude about alpha...

mathslover (mathslover):

I think, it is C) \(\alpha > \cfrac{1}{4} \)

Parth (parthkohli):

Yes, this is the easy part. But this is a more than one correct question.

ganeshie8 (ganeshie8):

Ohh

ganeshie8 (ganeshie8):

we need to find its upper bound also ?

Parth (parthkohli):

It may or may not have an upper bound. That's why the question is confusing.

Parth (parthkohli):

I kinda edited this question because it had a little hole. You may also want to think why a, b, c are distinct.

OpenStudy (ikram002p):

well for me instead of trying to reach the formula i can try to find which choice is correct

Parth (parthkohli):

lol

Parth (parthkohli):

Go ahead.

mathslover (mathslover):

Using AM-GM may give us something.. hmm!

ganeshie8 (ganeshie8):

Yup !

mathslover (mathslover):

lol Parth, can you do it from here? See, OpenStudy aims on providing the students a way to go through the question, we can not help you completely! You should try it first and let us know where you are getting problem!

mathslover (mathslover):

Haha! :)

Parth (parthkohli):

lulz

Parth (parthkohli):

Should I give it out now?

mathslover (mathslover):

Its up to you, we will not ask you to do so.. :P

ganeshie8 (ganeshie8):

clever way to escape typing the lengthy work @mathslover :P

hero (hero):

Just wait a sec

mathslover (mathslover):

lol , Thanks for the appreciation, it matters a lot for me , haha!

hero (hero):

x^2(a + b + c) + x(a + b + c) + 1(a + b + c) > 0 (a + b + c)(x^2 + x + 1) > 0 x^2 + x + 1 > 0

Parth (parthkohli):

\[a + b + c>0?\]

mathslover (mathslover):

@Hero - Yes, its right but what will it give to us? x^2 + x + 1 will always be positive for all real numbers

mathslover (mathslover):

And so, a + b + c > 0

Parth (parthkohli):

Right.

hero (hero):

I thought it would lead somewhere

mathslover (mathslover):

Yeah but I am still trying to get a nice looking equation from that ! Till now - No good results :(

OpenStudy (ikram002p):

humm so b,c also cant be any numbers :\ lolz good luck solving it !

ganeshie8 (ganeshie8):

Oka, I'm implementing the lazy @mathslover idea about AM-GM to complete the problem :) : a^2 + b^2 > 2ab b^2 + c^2 > 2bc c^2 + a^2 > 2ca ------------------ a^2 + b^2 + c^2 > ab+bc+ca (ab+bc+ca)/(a^2+b^2+c^2) < 1a

Parth (parthkohli):

YAY!

mathslover (mathslover):

lol @ganeshie8 , that's so kind of you to call me lazy...

OpenStudy (ikram002p):

i think all the choices cud be true according to c,b u can get out with

Parth (parthkohli):

Ganesh! Ganesh! Ganesh! Ganesh! Ganesh!

Parth (parthkohli):

<dingdingdingding> Alternative: The answer follows from an equality we learnt in the 9th grade.\[a^2 + b^2 + c^2 - ab - bc - ca = \dfrac{1}{2}\left(( a- b)^2 + (b - c)^2 + (c-a)^2\right)\]But since \(a,b,c\) are distinct, we have \(\dfrac{1}{2}\left(( a- b)^2 + (b - c)^2 + (c-a)^2\right) > 0\). In other words, \(a^2 + b^2 + c^2 - ab - bc - ca > 0\). So...

mathslover (mathslover):

Good work @ganeshie8 - My teachings worked finally :)

mathslover (mathslover):

So, it will be B) And C) both... !

Parth (parthkohli):

Yup.

mathslover (mathslover):

Okay, so, challenge solved by @ganeshie8 ! ~~~~~~~~~~Closed

Parth (parthkohli):

It was a nice time we had here, gentlemen.

mathslover (mathslover):

Gentlemen? Who? I can't see any "gentle" men here? (except me..)

Parth (parthkohli):

Funny.

mathslover (mathslover):

Let's see what Unkle has got there for us.